The solution set of pair of linear equation  3x+3y-3=0 and 5x+5y-5=0 

Grade:10

1 Answers

Shraddha Singh
2 years ago
Dear Student,
 
3x + 3y - 3 = 0
5x + 5y - 5 = 0
 
a1/a2 = 3/5
b1/b2 = 3/5
C1/c2 = -3/-5 = 3/5
 
Here, a1/a2 = b1/b2 = c1/c2 
Hence, the given equation has infinitely many solutions.
 
Hope it helps!
Thank you